# promptbox[1][] colframe=black, colback=white, sharp corners, boxrule=1pt, boxsep=8pt, left=0pt, right=0pt, top=0pt, bottom=0pt, #1

### v1

@@ 3.2.1. Selected data attributes
Our rationale for choosing these attributes is as follows. Although recommendation systems are often based on anonymized data (Resnick and Varian, 1997) and do not rely on profile pictures, names and profile pictures are still fundamental in personas (Grudin and Pruitt, 2002). They make personas real and relatable, serving as vital stimuli of user empathy (Dziobek et al., 2008). Furthermore, birth dates not only help establish the persona’s age, making them more recognizable to a certain age group, but also are pivotal for personalized content (Miaskiewicz and Kozar, 2011). However, due to ethical concerns, we omitted sensitive PII like phone numbers and Social Security Numbers.
@@
Demographic details, hobbies, and online interests play a crucial role in creating realistic personas. These attributes allow users to quickly connect with and relate to the personas through shared characteristics (Miaskiewicz and Kozar, 2011). Such connections foster user engagement and empathy, making interactions with personas more meaningful and relatable. Furthermore, online recommendation systems often utilize demographic data and personal preferences to tailor their offerings (Mohamed et al., 2019). This kind of personal data guides how online platforms categorize users and, subsequently, the type of content they receive. When users observe how demographic information and personal preferences impact the services or user experiences, they are more likely to disentangle the system’s opaqueness, understand how the system might use their privacy information, and consequently enhance their privacy awareness and literacy. Additionally, these data contribute significantly to the generation of longitudinal personal data. They influence a user’s weekly schedule by reflecting their lifestyle choices and priorities. These data also shape one’s browsing history and social media as individuals tend to browse and share content related to their hobbies and demographic identities.
@@
We incorporated details about devices and browsers into our personas, given their influence on online service personalization. Notably, certain studies, such as the one by Nikiforakis et al. (Nikiforakis et al., 2013), found that advertisers can use browser and device fingerprinting to tailor the ads for users. Hannak et al. (Hannak et al., 2014) observed that e-commerce platforms might offer different prices based on the user’s device or browser. Recognizing such distinctions can help users understand the link between their device/browser information and the system outcomes, which consequently enhances their privacy knowledge.
@@
Differing from traditional UX personas, we incorporated longitudinal personal data, like weekly schedules, location logs, browsing histories, and social media activities. From the user’s perspective, this richer dataset paints a more comprehensive picture of the persona’s life, fostering deeper user empathy. Previous work (Grudin and Pruitt, 2002) has also underscored the significance of longitudinal personal data in understanding personas. Rijn et al.’s study (Van Rijn et al., 2011) emphasizes the value of in-depth behavioral data in understanding personas. Thus, by observing these personas’ daily activities and interests, users can better empathize with them, seeing them as dynamic individuals with changing preferences.
From the system’s perspective, this temporal data is pivotal for online services to make contextual recommendations. Several studies have illustrated the use of social media (Camacho and Alves-Souza, 2018) and browsing histories (Tarus et al., 2017; Beigi et al., 2019; Rajendran and Sundarraj, 2021) in predicting user preferences. Consequently, this data not only amplifies user empathy but also impacts the tailored content they encounter. This insight helps users grasp the connection between their data and the content they receive, boosting their privacy awareness and possibly guiding their future privacy behaviors.
@@ 3.2.2. Data generation methodology
To create comprehensive persona data, we introduce a novel pipeline that augments the outputs of large language models using few-shot learning, contextualization, and chain-of-thoughts techniques. For readers’ reference, we have included detailed prompts and examples of few-shot learning in Appendix A.
Persona description: The foundational step in our process is generating a personal description, which informs subsequent data generation to ensure alignment. We utilize a template prompt coupled with few-shot learning (Brown et al., 2020), to guide GPT-4 in producing personally identifiable and demographic information. Users can customize the generation of their desired persona by providing other guidance as input.
Privacy attributes: We use GPT-4 and few-shot learning to parse the generated persona description and obtain attributes for each PII and demographic information to allow further modifications.
@@
Social media posts: To ensure realistic and consistent social media posts, we use the persona’s description and weekly schedule as contextual anchors in the prompt. Typically, social media posts might contain visuals, so we randomly add 0-2 images per post. If an image is integrated, we employ the “chain-of-thought” technique similar to how we generate profile pictures: the post content serves as input, generating a prompt for image synthesis. This prompt is then fed to the OpenAI DALL·E image generation API. This process enriches the realism of the persona, aligning visual content with the textual post. An illustrative prompt for social media posts is presented below.
{promptbox}
[title=Prompt for generating social media posts] Provide ideas for this person to write posts (limit the word to 140 words) based on the profile and location history: {profile} {location history}
Return a list of lists: ¡few-shot example posts¿
Output the posts in the following JSON format in plain text: { “time”: ¡time in string format¿, “address”: ¡address where this person shares the life¿, “content”: ¡content¿, “latitude”: ¡fake latitude¿, “longitude”: ¡fake longitude¿, “timezone”: ¡time zone¿, “locale”: ¡locale¿ }
{promptbox}
[title=Few-shot learning example for the generation of social media posts] “posts”: “[ [“2023-06-01 08:31:10”, “Starting my day with a delicious cup of coffee at my favorite coffee shop. Ready to conquer the world! #CoffeeLover #MorningMotivation”, “Coffee Shop - 123 Main Street, Brooklyn, New York 11207”], [“2023-06-01 18:00:34”, “Just got back from the grocery store. Stocked up on essentials for the week. #GroceryHaul #MealPrep”, “Grocery Store - 456 Broadway Avenue, Brooklyn, New York 11207”] ]”
{promptbox}
[title=Prompt for generating a prompt for generating an image associated with the post] Given the post {content}, output a descriptive prompt to generate a realistic life image, limit the word to 30 words:
@@ 3.2.3. Implementation details
To generate synthetic data, we used a Python script that interacted with the GPT-4 public API. We specified a maximum continuation length of $4,500$ tokens. Our approach to achieving few-shot learning involved utilizing the “FewShotPromptTemplate” available in the open-source Python library called “langchain222 https://pypi.org/project/langchain/”. Furthermore, we configured the GPT-4 model with a temperature parameter of $0.9$. The resulting images, generated using OpenAI DALL·E, were set to a size of $256\times 256$ pixels.
@@ 3.3. Personal data replacement
In the second stage, users can use the identities of the generated personas to interact with various online services. The sandbox replaces users’ demographic data within the Google account, real-time location, IP address, and web browsing history to match the persona’s attributes. When an online service requests this personal information, the system offers the synthetic data of the persona. For the service provider, this data seems genuine, allowing them to provide personalized content as though they were interacting with a real user. This approach offers users a risk-free platform to cognitively understand the tangible consequences of their privacy choices and emotionally empathize with the persona in a convincingly interactive environment without exposing their actual personal data. The subsequent sections detail the process of personal data replacement.
Google account: Since we choose online ads to represent system outcomes, data replacement for Google accounts primarily pertains to information within the Google Ad Center. Google Ad Center’s control portal [^3] https://myadcenter.google.com/controls allows users to customize the information provided to Google Ads, encompassing details of age, gender, language, relationship status, household income, education, industry, and homeownership. To substitute the profile data in the Google Ad Center with the privacy data of the persona, we create a Google account dedicated to the application. We use three open-source node.js libraries (“Puppeteer [^4] https://pptr.dev/”, “Puppeteer-extra”, and “Puppeteer-extra-plugin-stealth [^5] https://github.com/berstend/puppeteer-extra”) to automate the replacement of profile data. The replacement process consists of three steps: (1) After entering the personal profile page in Google Ad Center, we traverse the “aria-labels” of all elements of the page to identify the attributes that need to be replaced. (2) We extract the persona attributes from the database and process the data. (3) Then, we replace the values of target attributes with the persona’s corresponding information.
Geographical location: Personalized online advertisements are often tailored based on the user’s geographical location. Chrome browser supports location override. Replacement of geographical location involves two steps. First, based on the persona’s home address, we use the open-source mapping application OpenStreetMap’s geocoding API [^6] https://nominatim.openstreetmap.org/ui/search.html to obtain the latitude and longitude of the generated persona’s current address (based on their generated schedule). Then, we use the “setGeolocation” method from Puppeteer to modify the geographical location of the webpages based on the obtained latitude and longitude coordinates.
@@ 3.4.1. Privacy sandbox in action
Consider financial analyst Bob who wants to understand how private data impacts online ads. Using the Privacy Sandbox, he can generate a persona and act as the persona to browse websites that contain ads.
@@
    Providing Guidance for Persona Profile Generation: Bob chooses the “create a new persona” button and inputs his guidance to generate a persona that is similar to his profile. He enters “Financial analyst in Los Angeles, interested in online gaming and sports.” The “guidance” in this context acts as a seed or initial information. Users can provide as little or as much information as they feel comfortable with, ensuring flexibility while guarding their own private information. This information is not restricted to job titles or locations but could include hobbies, interests, or any other relevant information.
@@
    Initial Persona Profile Generation and Customization: Upon receiving the “guidance”, the Privacy Sandbox generates a preliminary persona for Bob. The generated persona for Bob is named Carlos Rodriguez, a 30-year-old Hispanic male living in Los Angeles. Carlos speaks both English and Spanish and has a bachelor’s degree in Business Administration. He works as a financial analyst and earns an annual income of \$75,000. He enjoys online gaming, watching YouTube videos, and checking sports scores. At this point, Bob can review the generated persona and modify any attributes of the persona, as shown in Fig. 4 (b).
@@
    Figure 5. Browsing Online Services with the Generated Persona: Users, after activating their generated persona, can interact with online services, observing the persona’s influence on targeted ads.
@@
    Browsing Online Services with the Generated Persona: Once Bob is satisfied with the persona, he can save this persona for future use. When he clicks on the “activate” button, the Privacy Sandbox uses a Chrome extension to replace Bob’s privacy data with the synthesized data of Carlos, the artificial persona. Bob’s privacy data, including his profile in the Google Ad Center, browsing history, real-time location, and IP address, are temporarily replaced. Bob can interact with websites and online services as usual, but online service providers will see him as Carlos and start providing him with personalized ads, customized content, and algorithmic decisions they would give to a person like Carlos. Bob can then experiment with different privacy settings (e.g., enabling/disabling the access to certain data for a website) and behaviors (e.g., visiting certain sites when the visits are tracked, voluntarily providing personal data to a service), experiencing how his user experience has changed as a result. For example, he may start to observe getting ads customized based on the attributes of Carlos.
@@ 3.4.2. Implementation
We developed the Privacy Sandbox with a React-based frontend and a backend powered by Flask and SQLite3. They communicate through HTTP requests for API access. The SQLite database stores different types of synthesized personal privacy data: persona profiles, schedules, browsing history, Twitter posts, and Facebook posts.
@@ 4.2.1. Study procedure
  • Phase 1: Quantitative evaluating generated personas Participants were presented with three personas on our developed privacy sandbox platform, one each randomly chosen from three distinct groups: personas generated with our approach, real personas, and personas directly generated using the GPT-4 model. We prepared eight personas for each group. The order of the personas was randomized. Participants were tasked to rate each persona’s clarity, completeness, credibility, consistency, and level of empathy using the five-point Likert scale.
@@
  • Phase 2: Qualitative investigation of generated personas To gain a deeper understanding of the user perceptions of generated personas and how real participants perceive different parts of them, we adopted a combined “Think Aloud” (Jääskeläinen, 2010) method and semi-structured interview approach. The experimenter first introduced the usage of the privacy sandbox to participants, ensuring that they understood how to generate and modify personas. Afterward, the selected generated personas were presented to the participants. As they navigated the personas, participants were instructed to vocalize their overall impressions and specifically comment on which elements of the persona’s privacy data enhanced or diminished the sense of realness. During the “Think Aloud” process, as participants shared their immediate feedback, researchers could interject with follow-up questions or ask for clarifications. If participants identified certain elements as inauthentic or felt adjustments were needed, the interface allowed them to directly modify the persona profiles, including attributes, avatars, weekly schedules, browsing history, and posts. They could either make direct modifications to the interface or verbally describe the desired changes. For every modification or suggestion, participants were asked to explain their reasoning. Each participant was exposed to two personas, counterbalanced, and selected from the eight personas. A detailed list of all the important attributes of these eight personas can be found in the appendix D.
@@
  • Phase 3: Analyzing Ad-Persona Connections In this phase, participants completed a task of correlating the persona information with the advertisements on given websites. The goal of this phase is to investigate whether users can perceive the correlation between privacy data and system outcome (e.g., target ads). Each participant completes the task for two personas that they have not seen before, randomly selected from the personas generated by our mechanism. For each persona, first, the participant read a persona using the privacy sandbox prototype. After reading, they clicked on a designated “active” button. This triggered the launch of a new browser window by the sandbox that automatically replaces their real personal data with the persona’s synthetic data. As explained in Section 3.3, the sandbox replaced persona attributes, browsing history, location, and IP address (as seen in Figure 5). Then, the participant was asked to read the home pages of two websites for each persona, randomly chosen from the five websites shown in Table 2. Participants were tasked to identify ads that are targeted to the current persona, record them in a spreadsheet, and explain how the ads relate to the persona in a think-aloud manner.

### v2

@@ 3.2.1. Selected data attributes
Our rationale for choosing these attributes is as follows. Although recommendation systems are often based on anonymized data (Resnick and Varian, 1997) and do not rely on profile pictures, names and profile pictures are still fundamental in personas (Grudin and Pruitt, 2002). They make personas real and relatable, serving as vital stimuli of user empathy (Dziobek et al., 2008). Furthermore, birth dates not only help establish the persona’s age, making them more recognizable to a certain age group, but are also pivotal for personalized content (Miaskiewicz and Kozar, 2011). However, due to ethical concerns, we omitted sensitive PII like phone numbers and Social Security Numbers.
@@
Demographic details, hobbies, and online interests play a crucial role in creating realistic personas. These attributes allow users to quickly connect with and relate to personas through shared characteristics (Miaskiewicz and Kozar, 2011). Such connections foster user engagement and empathy, making interactions with personas more meaningful and relatable. Furthermore, online recommendation systems often utilize demographic data and personal preferences to tailor their offerings (Mohamed et al., 2019). This kind of personal data guides how online platforms categorize users and, subsequently, the type of content they receive. When users observe how demographic information and personal preferences impact the services or user experiences, they are more likely to disentangle the system’s opaqueness, understand how the system might use their privacy information, and consequently enhance their privacy awareness and literacy. Additionally, these data contribute significantly to the generation of longitudinal personal data. They influence a user’s weekly schedule by reflecting their lifestyle choices and priorities. These data also shape one’s browsing history and social media posts, as individuals tend to browse and share content related to their hobbies and demographic identities.
@@
We incorporated details about devices, browsers, hobbies, and online interests into our personas, given their influence on online service personalization. Notably, certain studies, such as the one by Nikiforakis et al. (Nikiforakis et al., 2013), found that advertisers can use browser and device fingerprinting to tailor the ads for users. Hannak et al. (Hannak et al., 2014) observed that e-commerce platforms might offer different prices depending on the device or browser of the user. Furthermore, user hobbies and interests are often gathered through methods such as website tracking, online surveys, and social media monitoring, aiding in customer segmentation(Sari et al., 2016). Explicitly including such information in the persona allows the model to generate coherent person data (e.g., location history, browsing/search history, and social media posts) that reflect them. Therefore, when the user recognizes the system outcomes that relate to the hobbies and interests of the persona when interacting with the sandbox, they can understand their connection to the system’s access of personal data, contributing to the privacy knowledge of the user.
@@
Unlike traditional UX personas, we incorporate longitudinal personal data, such as weekly schedules, location logs, browsing histories, and social media activities. From the user’s perspective, this richer dataset paints a more comprehensive picture of the persona’s life, fostering deeper user empathy. Previous work (Grudin and Pruitt, 2002) has also underscored the significance of longitudinal personal data in understanding personas. Rijn et al.’s study (Van Rijn et al., 2011) emphasizes the value of in-depth behavioral data in understanding personas. Thus, by observing these personas’ daily activities and interests, users can better empathize with them, seeing them as dynamic individuals with changing preferences.
From the system’s perspective, these temporal data are pivotal for online services to make contextual recommendations. Several studies have illustrated the use of social media (Camacho and Alves-Souza, 2018) and browsing histories (Tarus et al., 2017; Beigi et al., 2019; Rajendran and Sundarraj, 2021) in predicting user preferences. Consequently, this data not only amplifies user empathy, but also impacts the tailored content they encounter. Such insights help users infer the connection between their data and the content they receive, contributing to empathy and privacy knowledge
@@ 3.2.2. Data generation methodology
To create comprehensive persona data, we introduce a novel pipeline that augments the outputs of large language models (LLMs) using few-shot learning (Brown et al., 2020), contextualization (Brown et al., 2020), and chain-of-thoughts techniques (Wei et al., 2022). Few-shot learning enables LLMs to quickly adapt to privacy persona generation using only a few examples. Contextualization allows LLMs to integrate additional context, such as weekly schedules, for more relevant output in tasks such as generating browsing history and social media posts. The chain-of-thoughts technique deconstructs complex tasks into simpler steps, enhancing the models’ ability to manage intricate privacy data generation. These three techniques collectively enhance LLMs’ performance by making them more adaptable, context-aware, and capable of generating complex data. For readers’ reference, we have included detailed prompts and examples of few-shot learning in Appendix A.
Persona description: The foundational step in our process is generating a personal description, which informs subsequent data generation to ensure alignment. We use a template prompt coupled with few-shot learning (Brown et al., 2020), to guide GPT-4 in producing personally identifiable and demographic information. Users can customize the generation of their desired persona by providing other guidance as input.
Privacy attributes: We use GPT-4 and few-shot learning to parse the persona description generated and obtain attributes for each PII and demographic information to allow further modifications.
@@
Social media posts: To ensure realistic and consistent social media posts, we use the persona’s description and weekly schedule as contextual anchors in the prompt. Typically, social media posts might contain visuals, so we randomly add 0-2 images per post. If an image is integrated, we employ the “chain-of-thought” technique similar to how we generate profile pictures: the content of the post serves as input, generating a prompt for image synthesis. This prompt is then fed to the OpenAI DALL-E image generation API. This process enriches the realism of the persona, aligning visual content with the textual content of the post. An illustrative prompt for social media posts is presented below.
Prompt for generating social media posts Provide ideas for this person to write posts (limit the word to 140 words) based on the profile and location history: {profile} {location history} Return a list of lists: ¡few-shot example posts¿ Output the posts in the following JSON format in plain text: { “time”: ¡time in string format¿, “address”: ¡address where this person shares the life¿, “content”: ¡content¿, “latitude”: ¡fake latitude¿, “longitude”: ¡fake longitude¿, “timezone”: ¡time zone¿, “locale”: ¡locale¿ }
Few-shot learning example for generating social media posts “posts”: “[ [“2023-06-01 08:31:10”, “Starting my day with a delicious cup of coffee at my favorite coffee shop. Ready to conquer the world! #CoffeeLover #MorningMotivation”, “Coffee Shop - 123 Main Street, Brooklyn, New York 11207”], [“2023-06-01 18:00:34”, “Just got back from the grocery store. Stocked up on essentials for the week. #GroceryHaul #MealPrep”, “Grocery Store - 456 Broadway Avenue, Brooklyn, New York 11207”] ]”
Prompt for generating a prompt for generating an image associated with the post Given the post {content}, output a descriptive prompt to generate a realistic life image, limit the word to 30 words:
@@ 3.2.3. Implementation details
To generate synthetic data, we used a Python script that interacted with the GPT-4 public API. We specified a maximum continuation length of $4,500$ tokens. Our approach to achieving few-shot learning involved utilizing the “FewShotPromptTemplate” available in the open-source Python library called “langchain222 https://pypi.org/project/langchain/”. Furthermore, we configured the GPT-4 model with a temperature parameter of $0.9$. The resulting images, generated using OpenAI DALL-E, were set to a size of $256\times 256$ pixels.
@@ 3.3. Personal data replacement
In the second stage, users can use the identities of the generated personas to interact with various online services. The sandbox replaces users’ demographic data within the Google account, real-time location, IP address, and web browsing history to match the persona’s attributes. When an online service requests this personal information, the system offers the synthetic data of the persona. For the service provider, these data seem genuine, allowing them to provide personalized content as if they were interacting with a real user. This approach offers users a risk-free platform to cognitively understand the tangible consequences of their privacy choices and emotionally empathize with the persona in a convincingly interactive environment without exposing their actual personal data. The subsequent sections detail the process of personal data replacement.
Google account: Since we choose online ads to represent system outcomes as explained in Section 3.1, data replacement for Google accounts primarily pertains to information within the Google Ad Center. Google Ad Center’s control portal [^3] https://myadcenter.google.com/controls allows users to customize the information provided to Google Ads, encompassing details of age, gender, language, relationship status, household income, education, industry, and homeownership. To substitute the profile data in the Google Ad Center with the privacy data of the persona, we create a Google account dedicated to the application. We use three open-source node.js libraries (“Puppeteer [^4] https://pptr.dev/”, “Puppeteer-extra”, and “Puppeteer-extra-plugin-stealth [^5] https://github.com/berstend/puppeteer-extra”) to automate the replacement of profile data. The replacement process consists of three steps: (1) After entering the personal profile page in Google Ad Center, we traverse the “aria-labels” of all elements of the page to identify the attributes that need to be replaced. (2) We extract the persona attributes from the database and process the data. (3) Then, we replace the values of target attributes with the persona’s corresponding information.
Geographical location: Personalized online advertisements are often tailored based on the user’s geographical location. Chrome browser supports location override. The replacement of the geographical location involves two steps. First, based on the persona’s home address, we use the open-source OpenStreetMap’s geocoding API [^6] https://nominatim.openstreetmap.org/ui/search.html to obtain the latitude and longitude of the generated persona’s current address (based on their generated schedule). Then, we use the “setGeolocation” method from Puppeteer to modify the geographical location of the webpages based on the obtained latitude and longitude coordinates.
@@ 3.4.1. Privacy sandbox in action
Consider Bob, a financial analyst who wants to understand how private data impacts online ads. Using the Privacy Sandbox, he can generate a persona and act as the persona to browse websites that contain ads.
@@
    Providing Guidance for Persona Profile Generation: Bob chooses the “create a new persona” button and enters his guidance to generate a persona that is similar to his profile. He enters “Financial analyst in Los Angeles, interested in online gaming and sports.” The “guidance” in this context acts as a seed or initial information. Users can provide as little or as much information as they feel comfortable with, ensuring flexibility while guarding their own private information. This information is not restricted to job titles or locations but could include hobbies, interests, or any other relevant information.
@@
    Initial Persona Profile Generation and Customization: Upon receiving the “guidance”, the Privacy Sandbox generates a preliminary persona for Bob. The persona generated for Bob is named Carlos Rodriguez, a 30-year-old Hispanic male living in Los Angeles. Carlos speaks both English and Spanish and has a bachelor’s degree in Business Administration. He works as a financial analyst and earns an annual income of \$75,000. He enjoys online gaming, watching YouTube videos, and checking sports scores. At this point, Bob can review the generated persona and modify any attributes of the persona, as shown in Fig. 4 (b).
@@
    Figure 5. Browsing Online Services with the Generated Persona: Users, after activating the Chrome extension using their generated persona, can interact with online services, observing the persona’s influence on targeted ads.
@@
    Browsing Online Services with the Generated Persona: Once Bob is satisfied with the persona, he can save this persona for future use. When he clicks on the “activate” button, the Privacy Sandbox will activate a Chrome extension that replaces Bob’s privacy data with the synthesized data of Carlos, the artificial persona. Bob’s privacy data, including his profile in the Google Ad Center, browsing history, real-time location, and IP address, are temporarily replaced. Bob can interact with websites and online services as usual, but online service providers will see him as Carlos and start providing him with personalized ads, customized content, and algorithmic decisions they would give to a person like Carlos. Bob can then experiment with different privacy settings (e.g., enabling/disabling the access to certain data for a website) and behaviors (e.g., visiting certain sites when the visits are tracked, voluntarily providing personal data to a service), experiencing how his user experience has changed as a result. For example, he may start to observe seeing ads customized based on the attributes of Carlos.
@@ 3.4.2. Implementation
We developed the Privacy Sandbox with a React-based fronte-nd and a back-end powered by Flask and SQLite3. They communicate through HTTP requests for API access. The SQLite database stores different types of synthesized personal privacy data: persona profiles, schedules, browsing history, Twitter posts, and Facebook posts.
@@ 4.2.1. Study procedure
  • Phase 1: Quantitative evaluation of generated personas Participants were presented with three personas on our developed privacy sandbox platform, one each randomly chosen from three distinct groups: personas generated with our approach, real personas, and personas directly generated using the GPT-4 model (detailed in Section 4.2.2). Each group contains eight personas. The order of the personas was randomized. Participants were tasked to rate each persona’s clarity, completeness, credibility, consistency, and level of empathy using the five-point Likert scale.
@@
  • Phase 2: Qualitative investigation of generated personas To gain a deeper understanding of user perceptions of generated personas and how real participants perceive different parts of them, we adopted a combined “Think Aloud” (Jääskeläinen, 2010) method and semi-structured interview approach. The experimenter first introduced the use of the privacy sandbox to the participants, ensuring that they understood how to generate and modify personas. Afterward, the selected generated personas were presented to the participants. As they navigated the personas, participants were instructed to vocalize their overall impressions and specifically comment on the elements of the persona’s privacy data (as detailed in Section 3.2.1 and Figure 4) enhanced or diminished the sense of realness. During the “Think Aloud” process, as participants shared their immediate feedback, researchers could interject with follow-up questions or ask for clarifications. If participants identified certain elements as inauthentic or felt adjustments were needed, the interface allowed them to directly modify the persona profiles, including attributes, avatars, weekly schedules, browsing history, and posts. They could either make direct modifications to the interface or verbally describe the desired changes. For every modification or suggestion, participants were asked to explain their reasoning. Each participant was exposed to two personas, counterbalanced, and selected from the eight personas. A detailed list of all the important attributes of these eight personas can be found in the Appendix D.
@@
  • Phase 3: Analyzing ad-persona connections In this phase, participants completed the task of correlating the persona information with the advertisements on given websites. The goal of this phase is to investigate whether users can perceive the correlation between privacy data and system outcome (e.g., target ads). Each participant completes the task for two personas that they have not seen before, randomly selected from the personas generated by our mechanism. For each persona, first, the participant read a persona using the privacy sandbox prototype. After reading, they clicked on a designated “active” button. This triggered the launch of a new browser window by the sandbox that automatically replaces their real personal data with the persona’s synthetic data. As explained in Section 3.3, the sandbox replaced persona attributes, browsing history, location, and IP address (as seen in Figure 5). Then, the participant was asked to read the home pages of two websites for each persona, randomly chosen from the five websites shown in Table 2. Participants were tasked to identify ads that are targeted to the current persona, record them in a spreadsheet, and explain how the ads relate to the persona in a think-aloud manner.